OpenStudy (anonymous):

3x+2y=4 x+3y=-1 solve using addition

HanAkoSolo (jamierox4ev3r):

are you sure you use addition? don't you mean substitution or elimination?

HanAkoSolo (jamierox4ev3r):

@amyamanda

OpenStudy (anonymous):

it says addition

HanAkoSolo (jamierox4ev3r):

then just add the two systems together. what is 3x+x?

OpenStudy (anonymous):

4x

HanAkoSolo (jamierox4ev3r):

and 2y+3y? Good!, btw

OpenStudy (anonymous):

5y

HanAkoSolo (jamierox4ev3r):

then 4+(-1)?

OpenStudy (anonymous):

3

HanAkoSolo (jamierox4ev3r):

right so put it all together and you get... :)

OpenStudy (anonymous):

so 4x+5y=3

HanAkoSolo (jamierox4ev3r):

right! good job :)

OpenStudy (anonymous):

r u sure that's right, that doesn't seem right

HanAkoSolo (jamierox4ev3r):

why? is it not a choice, because as you solve by addition, addition means to add, so this should be your answer

HanAkoSolo (jamierox4ev3r):

unless i have misunderstood your question...

OpenStudy (mathmate):

"Addition" does not stop you from multiplying by a + or - factor before you add. Essentially it would be elimination.

OpenStudy (anonymous):

? I'm confused

HanAkoSolo (jamierox4ev3r):

so you would use elimination? hmmm i must have learned it differently

OpenStudy (anonymous):

@mathmate does that mean that 4x+5y=3 is incorrect?

OpenStudy (mathmate):

As far as I know, I would not accept 4x+5y=3 as the solution to a system of equations.

OpenStudy (anonymous):

ok @mathmate could you please break down step by step for me

HanAkoSolo (jamierox4ev3r):

right i was just confused when she said solve by addition. However for this case, substitution seems more applicable

HanAkoSolo (jamierox4ev3r):

or is it just me? idk

OpenStudy (mathmate):

The solution set to a system of two equations would typically (but not always) be S=(x,y) where x,y are numbers.

OpenStudy (mathmate):

Yes, substitution and elimination are both feasible.

OpenStudy (mathmate):

However, elimination is a kind of addition, so that's what I suggest, not because it is simpler.

OpenStudy (anonymous):

ok how do I use elimination?

OpenStudy (mathmate):

Elimination means that you multiply one or both equations with certain numbers so that the coefficients of a variable are identical. For example, if you multiply the second equation by -3, you will get 3x+2y=4 -3x-9y=3 Now you need to add the two equations and solve for y because the x's will cancel out when you ADD.

OpenStudy (anonymous):

y=-1

OpenStudy (anonymous):

@mathmate y=-1

OpenStudy (mathmate):

Very good, now all you have to do is to put y=-1 into one of the original equations and solve for x.

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

independent equation because it only has one solution

OpenStudy (mathmate):

Good observation! Now the last step (whenever you solve a system of equations) is to put back the solution set into the original equations and see if they "work". What is your solution set?

OpenStudy (anonymous):

6-2=4 2-3=-1

OpenStudy (mathmate):

Perfect, they match. The solution set is S=(2,-1) i.e. x=2, y=-1, is that ok?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ty!

OpenStudy (mathmate):

yw!